If a pitcher throws a fastball at 40m/s. If there is a 12m/s wind blowing from first base towards third base, how fast does the batter see the ball travel? How much time does the batter have to hit the ball if the distance from the pitcher’s mound to home plate is 60.5 feet (18.4 meters)?
1
Expert's answer
2020-03-05T09:37:21-0500
"u=v-v_w=40-12=28\frac{m}{s}"

"t=\frac{d}{u}=\frac{18.4}{28}=0.657\ s"